OpenStudy (anonymous):

How to find paramatric equation for a given curve? Anyone who can help me finding the paramatric equation of the curve x^(2/3) + y^(2/3) = a^(2/3)

hartnn (hartnn):

to find the parametric equation of the curve, u need to find x any y separately as the function of a....x=f(a),y=f(a).....such that the above equation satisfies. for this equation,we use the fact that sint^2 + cost ^2 =1 so my substitution will be x=a (cost)^3 y=a(sint)^3 (because i saw the power of 2/3 in x,when i needed 2) if u put these ,u will get RHS as a^2/3 so the parametric equations are x=a (cost)^3 y=a (sint)^3

hartnn (hartnn):

got it @vishwajeet

OpenStudy (anonymous):

@hartnn thanks! Would you please do a favour of doing the same for a parabola and a hyperbola!

hartnn (hartnn):

there are various equations for parabola and hyperbola the parametric substitution will depend on the equation... if u give an equation,then i can tell u specific substitution for it...

OpenStudy (anonymous):

@hartnn y^2 = 4x +5 and (x^2)/16 - (y^2)/9 = 1

hartnn (hartnn):

for parabola, u can put x=t^2-5/4,(so that 4x becomes,4t-5...to eliminate constant 5).,then u get y as 2t(y^2=4*t^2)..... for hyperbola,are u familiar with hyperbolic functions sinh x,cosh x?